# Personalizing story generation by inferring author styles

### v1

@@ 7.2 Story Understanding and Generation
Prior work on persona-aware story generation (Yunusov et al., 2024; Bae and Kim, 2024; Zhang et al., 2022; Chandu et al., 2019) defines personas using discrete attributes like personality traits, demographics, or hobbies. Similarly, (Zhu et al., 2023) explore story style transfer across pre-defined domains (e.g., fairy tales, martial arts, Shakespearean plays). In contrast, we mimic an individual author’s writing style based on their history. Our approach differs by (1) inferring long-form author personas—descriptions of an author’s style from their past works, rather than relying on demographics, and (2) handling long-form story generation, averaging 1500 tokens per output, exceeding typical story lengths in prior research.
8 Conclusion
We introduce the task of personalizing authors’ story-writing styles across five sources with a two-stage approach that infers implicit characteristics as an Author Writing Sheet to guide LLMs via persona descriptions and story rules. Evaluations show that summarizing an author’s history via a Writing Sheet or Summary improves faithfulness to writing history and similarity to the ground truth. Personalization is more effective for sources with diverse writing prompts and distinctive author traits (e.g., Reddit, New Yorker) and for narrative categories like Creativity and Language Use, which are less writing prompt-dependent.
@@ 9 Limitations
Limited Author History Size:
Table 4 (70% of the Stories/Author column) shows that N.Magazine and New Yorker authors have an average of 3 stories in their profiling set, whereas Reddit and AO3 authors have around 10. A larger profiling set likely enables more accurate inference of an author’s story-writing characteristics, leading to improved personalization. However, the cognitive demands of writing long stories result in fewer stories per author available online, thereby limiting the scalability of personalized story generation research.
@@
We validate the precision of our Author Writing Sheet through human evaluation, ensuring that the Claim-Evidence pairs accurately describe the author’s story-writing style. However, we do not assess recall, i.e., whether all of the author’s story-writing characteristics are captured. Measuring recall is challenging due to the inherent subjectivity of the task, making it impractical to ask humans to independently construct Writing Sheets for comparison.
@@
Our work explores role-playing for personalized story generation by prompting an LLM with tailored persona descriptions and story rules. Future research can extend this by incorporating multi-agent systems, where specialized agents focus on different narrative categories to enhance personalization (Huot et al., 2024; Bae and Kim, 2024). Additionally, fine-tuning LLMs to explicitly reason over author history could be investigated as a method for better aligning generated content with individual author preferences (Salemi et al., 2025; Shashidhar et al., 2024; Shaikh et al., 2024).
Automatic Evaluation Methods:
Evaluating long-form text remains challenging for LLMs, particularly for creative outputs that involve nuanced aspects like subtext, which are difficult to assess automatically and often require human evaluation (Subbiah et al., 2024a).
@@ 10 Ethical Considerations
All data used in this study were manually collected from publicly available sources [^4] Permission was obtained from the authors of the Storium dataset for its inclusion in Mythos., adhering to the data usage and crawling policies of the respective websites. To comply with copyright constraints, we will release only the links to the stories used in this research, rather than the stories themselves, following the approach of (Chakrabarty et al., 2024).
Fully anonymizing our dataset is not feasible [^5] Storium authors already anonymize their dataset. as the research involves mimicking specific authors’ writing behaviors. We note that even if the usernames are anonymized, it is still possible to infer an author’s publicly available identity from the story link. Therefore, for each author for every source, we release their usernames along with corresponding writing prompts, and author-written stories in the profiling and generation set of our dataset. However, the dataset remains limited in scale, and no additional metadata, including demographic information, is included, thereby reducing potential risks. To further mitigate unforeseen harms, the dataset will be released under an **Educational or Academic Research, Non-Commercial Use** license, following (Akoury et al., 2020).
@@ Appendix A Dataset
In this section, we describe the details of the construction of our dataset which includes five distinct story-writing sources, Reddit [^6] https://www.reddit.com/r/WritingPrompts/, AO3777https://archiveofourown.org/, Storium [^8] https://storium.cs.umass.edu/, N.Magazine [^9] https://www.narrativemagazine.com/, and New Yorker [^10] https://www.newyorker.com/.
@@
To ensure high-quality and diverse content, we apply several constraints during dataset collection. Stories are limited to a length of 500 to 1500 words. To avoid contamination by LLM pretraining data, only recent stories published after November 2023 are included for Reddit and AO3 (Zhou et al., 2023; Magar and Schwartz, 2022). Not-Safe-For-Work (NSFW) and explicit content is excluded using automatic tagging and manual verification. Additionally, we manually inspect all stories to remove elements that may reveal author identities, such as URLs and links to their public profiles on websites.
@@
The dataset contains 590 stories from 64 authors, with an average story length of 1500 tokens. Detailed statistics are provided in Table 4. As shown in Table 1, our dataset uniquely combines diverse story-writing settings and provides connections that link stories written by the same author, distinguishing it from existing story-writing datasets that either lack Author IDs or do not establish such links.
@@
Figure 4 shows the plots depicting the percentage-normalized frequency of the top 10 themes for each source. To derive these themes, we prompt GPT-4o with each writing prompt to generate up to five one-word themes. We then aggregate the themes across all writing prompts and compute the normalized frequency by dividing the count of each theme by the total number of themes for that source.
Overall, we observe that Reddit (Figure 4(a)), AO3 (Figure 4(b)), and Storium (Figure 4(e)) contain a higher prevalence of “conflict” themes, whereas N.Magazine (Figure 4(c)) and New Yorker (Figure 4(d)) focus more on family, loss, and identity. This distinction reflects the nature of story-writing settings across these sources: amateur stories tend to emphasize direct conflict resolution in their prompts, while expert-written stories explore deeper aspects of human relationships.
Among Reddit, AO3, and Storium, Reddit exhibits a flatter distribution with greater thematic diversity, spanning identity, power, survival, humor, adventure, magic, and the supernatural. In contrast, AO3 has a sharper distribution, dominated by recurring themes of relationships, friendships, love, family, and secrets, reflecting a more domestic story-writing setting. Storium also shows a sharper distribution but favors themes such as space, danger, leadership, and war, highlighting a preference for war-centric narratives in collaborative storytelling.
@@
N.Magazine 32 18 14 10 3 1745
NewYorker 38 22 16 10 4 1950
Overall 590 383 207 64 8 1517
(a) Reddit
(b) AO3
(c) N.Magazine
(d) New Yorker
(e) Storium
Figure 4: Source-wise Percentage Normalized Frequency of the top 10 themes.
@@ 4 $A_{0}\leftarrow\emptyset$
Algorithm 1 Constructing the Author Writing Sheet from a profiling set of stories
@@ B.1 Cost Analysis of Author Writing Sheet
In this section, we analyze the cost of generating the Author Writing Sheet (see Section 3) and compare it with the cost of generating the Author Writing Summary (see Writing Summary) as a function of the profiling set size ($|P|$) using GPT-4o. We calculate the average cost per sample (a writing prompt and its corresponding story) as 0.06\$ for the Author Writing Sheet and 0.02\$ for the Author Writing Summary considering 2.50\$ per 1M tokens for input, and 10.00\$ per 1M tokens for output [^12] https://openai.com/api/pricing/.
@@ D.1 GPT4o-as-a-judge evaluation
Prompt for Evaluating AI-Generated Stories Against Human-Written References
@@
Figure 20 shows the prompt for GPT4o-as-a-judge evaluation for Faithfulness to Writing History using the Author Writing Summary as the reference. Figure 21 shows the for GPT4o-as-a-judge evaluation for Similarity to Author Story. We prompt GPT4o separately with each narrative category and its description to obtain the winner among the pair of stories (Saha et al., 2024).
@@ D.2 Traditional Metrics
We complement GPT-4o-as-a-judge evaluations with traditional metrics (Ivey et al., 2024; Huot et al., 2024) across three criteria: (1) lexical overlap, measured using BLEU and ROUGE-L (Xie et al., 2023); (2) story diversity, assessed via homogenization scores (Padmakumar and He, 2023); and (3) stylistic similarity, evaluated separately for the author history and the ground-truth author story using LUAR (Rivera-Soto et al., 2021). Specifically, we compute two stylistic similarity scores: one measuring the cosine similarity between the LUAR embeddings of the generated story and the mean embeddings of the stories in the author’s profiling set (author history), and another comparing the generated story to the ground-truth author story. Additionally, we report generated story lengths in tokens.
@@ E.1 Faithfulness to Writing History
Claims from the Author Writing Summary for Language Use
1. A conversational and informal tone is employed, using humor and sarcasm.
2. Dialogue reflects characters’ distinct voices and contributes to the comedic tone.
3. Vivid and sometimes crude language adds to the immersive nature of the stories.
@@
"At the Library at the End of the Universe, where dawn and dusk are the only times that exist, the Librarians are tasked with venturing into Bookworlds to fix stories that have gone astray—but what happens when the Chronicle itself begins to rewrite *their* story?"
Oracle Story
In the grand cosmic tapestry, where all threads converge into the singular, glittering endpoint known as the Library at the End of the Universe, there lies a place so vast, so staggeringly infinite, that it defies all comprehension. You just won’t believe how vastly, hugely, mind-bogglingly big it was. Imagine a labyrinth of endless corridors, each lined with shelves that stretch beyond the horizon, filled with books of every conceivable and inconceivable nature. In this whimsical realm, the concept of time is a mere suggestion, a faint whisper of dawn and dusk that play tag across the vaulted ceilings, casting elongated shadows that danced like giddy specters…….With books in hand and hearts alight with curiosity, they stepped forward into the unknown, ready to write the next chapter in the endless, wondrous saga of the Library at the End of the Universe.
Writing Summary Story
In the Library at the End of the Universe, …. But today, chaos decided to go on a caffeine-fueled rampage. The dawn had decided to join forces with dusk, creating a bizarre twilight that left the Librarians squinting like moles in a disco…Amongst the squad of book-botherers, there was Reginald, a Librarian so obsessed with order that he arranged his socks by the Dewey Decimal System. …Gertie grinned, her parrot squawking yet another Shakespearean soliloquy. Äye, who knew we’d end up in a romance novel, eh? But hey, at least we’re not in a tragedy!
Average Author Story
In the Library at the End of the Universe, where dawn and dusk were the only times that existed, a peculiar light filtered through the stained glass windows, casting kaleidoscopic shadows across endless shelves of books…..Livia, a seasoned Librarian with silver streaks in her auburn hair, stood before the Chronicle, her heart pounding with an uneasy rhythm…. "This isn’t right," she murmured, her voice swallowed by the cavernous room. "The Chronicle should not change on its own." Beside her, Eamon, a younger Librarian with keen eyes and an inquisitive mind, frowned. "Perhaps it’s a reflection of a story gone wrong? One of the Bookworlds may be bleeding into our reality."
Table 8: Example illustrating a case where the Oracle story deviates from the author’s documented story-writing characteristics in the Language Use category of the Author Writing Summary for a Storium author. Highlighted text in the Average Author Story shows partial alignment with the Author Writing Summary, while highlighted text in the Writing Summary story demonstrates stronger alignment. In contrast, the Oracle story lacks alignment, leading to the Average Author Story being preferred over the Oracle story and the Writing Summary story being preferred over the Average Author story. This explains the lower win-rate of the Oracle method compared to the Writing Summary method.
Table 8 shows an example where the Oracle story deviates from the author’s documented story-writing characteristics as described in the Author Writing Summary for a Storium author. The author’s profiling set consists of three stories that consistently employ informal and colloquial language using humor and sarcasm, often conveyed through dialogue. However, for the new writing prompt in the generation set, the author’s style shifts to a more serious and mystic tone, diverging from their previously established preferences. In contrast, the Average Author method partially aligns with the Author Writing Summary by incorporating dialogue, while the Writing Summary method exhibits stronger alignment by more accurately capturing the author’s use of informal language and comedic tone. The Oracle story, however, fails to align with the author’s established style. This phenomenon results in the Average Author story being preferred over Oracle, and the Writing Summary story being preferred over the Average Author story, explaining the lower win-rate of the Oracle method compared to the Writing Summary method in GPT-4o-as-a-judge evaluation for Faithfulness to Author History.
E.2 Effect of Persona on Similarity to Author Story

### v2

@@ 8.2 Story Understanding and Generation
Prior work on persona-aware story generation (Yunusov et al., 2024; Zhang et al., 2022; Chandu et al., 2019) defines personas using discrete attributes like personality traits, demographics, or hobbies. Similarly, (Zhu et al., 2023) explore style transfer across predefined domains (e.g., fairy tales, martial arts, Shakespearean plays). In contrast, we mimic an individual author’s style based on their history. Our approach differs by (1) inferring long-form author personas descriptions of style from past works, rather than relying on demographics, and (2) handling long-form story generation, averaging 1500 tokens per output, beyond typical lengths in prior work.
9 Conclusions and Future Work
In this paper, we introduced the task of personalizing authors’ story-writing styles across five sources by proposing a new dataset, Mythos, containing 3.6k stories. We also proposed a two-stage approach that infers implicit characteristics as a high-quality Author Writing Sheet, to guide LLMs via persona descriptions and story rules. Extensive automated and human evaluation showed that our method outperforms baselines for both faithfulness to the author’s history and similarity to the ground-truth author story. However, our methods still significantly lagged behind an Oracle method, by as much as 24% in some ways, which highlights the difficulty of the task and the need for future work. We also found that personalization is more effective for sources with diverse writing prompts and distinctive author traits (e.g., Reddit, New Yorker), and for narrative categories like Creativity and Language Use, which are less prompt-dependent.
Future work can address the limitations of the Author Writing Sheet by improving its coverage of relevant stylistic information despite context length constraints, and by retrieving or re-ranking information relevant to the given writing prompt during generation. Another promising direction is to explore multi-agent systems for story-writing personalization, with specialized agents handling different narrative categories to improve the extraction and application of author stylistic traits especially for categories like Plot and Development that are hard to personalize.
@@ 10 Limitations
Limited size of Profiling Set for some sources:
Table 6 shows that when using 70% of each author’s available stories for profiling (from Stories/Author column), authors from Narrative Magazine and New Yorker have only about 2 stories on average, compared to 5 for Storium and around 10 for AO3. Reddit is the exception, with authors having an average of 38 stories in their profiling set. A larger profiling set likely enables more accurate inference of an author’s story-writing characteristics, leading to better personalization. However, our data collection is constrained by practical limitations. For all sources except Reddit, we manually collect publicly available stories without web scraping, which restricts the number of samples we can obtain. Furthermore, for expert sources like Narrative Magazine and New Yorker, fewer stories per author are available online, limiting the scalability of personalized story generation. For Reddit, we adhere to Reddit’s data collection policies and use the official API for dataset construction.
@@
Our method for generating the Author Writing Sheet is limited in its ability to capture all of an author’s story-writing characteristics. This limitation arises from practical context length constraints and the challenges of long-context reasoning during the iterative merging step, which forces us to restrict the Author Writing Sheet to 10 characteristics per narrative category. Regarding human evaluation, we validate only the precision of the Author Writing Sheet—ensuring that the extracted Claim-Evidence pairs accurately reflect the author’s style—but do not assess recall, i.e., whether all relevant characteristics are captured. Measuring recall is challenging due to the inherent subjectivity of the task, making it impractical to ask annotators to independently construct Writing Sheets for comparison.
@@
Our work explores role-playing for personalized story generation by prompting an LLM with tailored persona descriptions and story rules. Future research can extend this by incorporating multi-agent systems, where specialized agents focus on different narrative categories to enhance personalization (Huot et al., 2024; Bae and Kim, 2024), especially for categories like Plot and Development, which are less conducive to personalization. Additionally, while our experiments show that fine-tuning improves personalization over prompting for LLaMA 3.1 8B (Section F.2 and Table 21), our work sets a precedent for using fine-tuning to better align LLMs with individual author preferences. Future work can build on this by incorporating explicit reasoning over author history to further improve personalization (Salemi et al., 2025; Shashidhar et al., 2024; Shaikh et al., 2024).
Subjectivity in Evaluation:
Evaluating long-form creative text remains challenging for LLMs due to the involvement of nuanced aspects like subtext (Subbiah et al., 2024a). Despite this subjectivity, we obtain fair Cohen’s Kappa agreement for both human-human and human-model evaluations. Future work can explore improving evaluation methods by using aspect-based checklists that capture finer-grained narrative details (Lee et al., 2025).
@@ 11 Ethical Considerations
Dataset Collection:
All data used in this study were manually collected from publicly available sources, [^5] Permission was obtained from the authors of the Storium dataset for its inclusion in Mythos. adhering to the data usage and crawling policies of the respective websites. For Reddit, we collected data exclusively using the official Reddit API, following Reddit’s data access policies. [^6] https://www.redditinc.com/policies/data-api-terms To comply with copyright constraints, we will release only the URLs linking to the original stories, rather than the story texts themselves, following the approach of (Chakrabarty et al., 2024).
Data Anonymization and Licensing:
Fully anonymizing our dataset is not feasible, [^7] Storium authors already anonymize their dataset. as the research involves mimicking specific authors’ story writing styles. Even if usernames were anonymized, authors’ publicly available identities could still be inferred through the story links. Therefore, for each author and source, we release their usernames along with the corresponding writing prompts and author-written stories in the profiling and generation sets. No additional metadata, including demographic information, is included, minimizing potential privacy risks. To further mitigate unforeseen harms, the dataset will be released under an **Educational or Academic Research, Non-Commercial Use** license, following (Akoury et al., 2020).
@@ Appendix A Dataset
In this section, we describe the details of the construction of our dataset which includes five distinct story-writing sources, Reddit [^8] https://www.reddit.com/r/WritingPrompts/, AO3999https://archiveofourown.org/, Storium [^10] https://storium.cs.umass.edu/, Narrative Magazine [^11] https://www.narrativemagazine.com/, and New Yorker [^12] https://www.newyorker.com/.
@@
To ensure high-quality and diverse content, we apply several constraints during dataset collection. Stories are limited to a length of 500 to 1500 words. To avoid contamination by LLM pretraining data, only recent stories published after the release of ChatGPT [^13] https://help.openai.com/en/articles/6825453-chatgpt-release-notes November 2022 are included for Reddit and AO3 (Zhou et al., 2023; Magar and Schwartz, 2022). Not-Safe-For-Work (NSFW) and explicit content is excluded using automatic tagging and manual verification. Additionally, we manually inspect all stories to remove elements that may reveal author identities, such as URLs and links to their public profiles on websites.
@@
The dataset contains 3.6k stories from 112 authors, with an average story length of 1500 tokens. Reddit forms the largest subset of our dataset with 3.2k stories given the larger availability of stories on the platform and the ease of data selection using their API  [^14] https://www.reddit.com/r/reddit/comments/145bram/addressing_the_community_about_changes_to_our_api/. Detailed statistics are provided in Table 6. As shown in Table 1, our dataset uniquely combines diverse story-writing settings and provides connections that link stories written by the same author, distinguishing it from existing story-writing datasets that either lack Author IDs or do not establish such links.
@@
Figure 3: Story-writing themes for the Reddit source covered in Mythos.
The pie chart Figure 3 shows the distribution of topic keywords (story-writing themes) obtained from the Reddit source of Mythos using BERTopic Grootendorst (2022) on the writing prompts. We see that our dataset covers a wide range of themes including Magic, Detectives, and Aliens exploring various genres of story-writing.
@@
N.Mag 32 18 14 10 3 1745
N.York 38 22 16 10 4 1950
Overall 3682 2525 1157 112 16 1517
@@ 4 $A_{0}\leftarrow\emptyset$
Algorithm 1 Constructing the Author Writing Sheet from the profiling set
@@ B.1 Cost Analysis of Author Writing Sheet
In this section, we analyze the cost of generating the Author Writing Sheet (see Section 3) and compare it with the cost of generating the Author Writing Summary (see Summary) as a function of the profiling set size ($|P|$) using GPT-4o. We calculate the average cost per sample (a writing prompt and its corresponding story) as 0.06\$ for the Author Writing Sheet and 0.02\$ for the Author Writing Summary considering 2.50\$ per 1M tokens for input, and 10.00\$ per 1M tokens for output [^16] https://openai.com/api/pricing/.
@@ D.1 LLM-as-a-judge evaluation
Prompt for Evaluating Similarity to Author Story
@@
Figure 20 shows the prompt for LLM-as-a-judge evaluation for Faithfulness to Writing History using the Author Writing Summary as the reference. Figure 21 shows the for LLM-as-a-judge evaluation for Similarity to Author Story. We prompt OpenAI o4-mini separately with each narrative category and its description to obtain the winner among the pair of stories (Saha et al., 2024).
@@ D.2 Traditional Metrics
We complement LLM-as-a-judge evaluations with traditional metrics (Ivey et al., 2024; Huot et al., 2024) across three criteria: (1) lexical overlap, measured using BLEU and ROUGE-L (Xie et al., 2023); (2) story diversity, assessed via homogenization scores (Padmakumar and He, 2023); and (3) stylistic similarity, evaluated separately for the author history and the ground-truth author story using LUAR (Rivera-Soto et al., 2021). Specifically, we compute two stylistic similarity scores: one measuring the cosine similarity between the LUAR embeddings of the generated story and the mean embeddings of the stories in the author’s profiling set (author history), and another comparing the generated story to the ground-truth author story. Additionally, we report generated story lengths in tokens.
@@ E.1 Faithfulness to Writing History
E.1.1 Sheet and Summ achieve best faithfulness to writing history:
@@
"Welcome to Hell. To leave, all you must do is finish seven impossible tasks decided by your seven greatest sins. Four people have made it out so far. Will you be the fifth?"
Author Writing Summary (Language Use)
The author frequently uses dialogue to reveal character traits and dynamics. The author often employs vivid imagery to create an immersive atmosphere. The author often uses rhetorical questions to convey a character’s internal conflict or uncertainty.
Average Author
The air was thick with an acrid, sulfurous scent as Lucy stood at the threshold of Hell, a place she’d never believed in until now. A figure emerged from the shadows, tall and imposing, yet with a surprisingly warm smile. "Welcome to Hell," he said, his voice smooth and oddly comforting. "To leave, all you must do is finish seven impossible tasks decided by your seven greatest sins. Four people have made it out so far. Will you be the fifth?"
Sheet
I awoke on a bed of ash and cinders, the air around me thick with the scent of sulfur and regret. A laugh, echoing and absurd, filled the air—a demon guide, amused by my struggle, appeared at my side.
"All this for a toy soldier?" it chortled, eyes gleaming with mischief. "You must learn to let go, or you’ll never find the exit."
As I navigated the shifting corridors, I pondered the truth in its words. Each step forward was a step away from the past, from the child clutching stolen dreams.
Delta
Flames licked the edges of a crimson sky, and the ground beneath my feet felt alive, pulsing with a malevolent energy. I stood on the precipice of an endless abyss, the cacophony of wailing souls echoing in my ears.
Emerging from the maze, I found myself in a grand hall, gilded and opulent, dripping with wealth beyond imagination. Coins rained from the ceiling, and jewels sparkled like fallen stars.
Greed beckoned me to take, to hoard, to drown in the riches that lay at my feet. But I knew the price of succumbing; this treasure would bury me alive.
Table 10: Example where Sheet outperforms Delta for Faithfulness to Author History. Sheet reflects the Author Writing Summary through vivid imagery, character-driven dialogue, and rhetorical questions, whereas Delta emphasizes setting and symbolism but lacks interactive dialogue and internal reflection.
Table 10 shows an example for a Reddit author where the Sheet method outperforms Delta in terms of Faithfulness to Author History for the Language Use category. The Author Writing Summary highlights stylistic traits such as vivid imagery, dialogue, and rhetorical questions. The Sheet captures these through immersive narration and reflective exchanges with a demon guide. In contrast, Delta focuses on abstract setting descriptions and symbolic temptation but omits interactive dialogue and rhetorical structure, making it less faithful to the author’s characteristic style.
E.1.2 Persona descriptions benefit amateur sources more than expert sources for Summ
Author Writing Summary for Language Use
The author employs conversational and accessible language to engage the reader. The author uses humor and wit in dialogue to add depth to characters and situations. The author frequently uses rhetorical questions to engage the reader and provoke thought. The author uses repetition for emphasis and to reinforce key themes or ideas.
Persona Description for Language Use
Your language use is conversational and accessible, engaging readers with humor and wit. You employ vivid imagery to create immersive settings and use rhetorical questions to provoke thought. Repetition reinforces key themes, while a mix of short and long sentences controls pacing and builds tension.
Writing Prompt
The reason humanity has never experienced any world-ending events is that future humans keep travelling back in time to stop them. You know this because there’s one of them in front of you now, telling you that you must absolutely not go into work today.
Summ
Evelyn’s mornings were a symphony of simplicity… "Whoa, what year is it?" he blurted out, eyes wide with what could only be described as temporal jet lag… "Listen, Evelyn, you absolutely cannot go to work today." Evelyn blinked, her skepticism as thick as the cream in her coffee. "And why is that? Did the universe finally decide to give me a spontaneous day off?"… Evelyn took a slow sip of her coffee, considering her options. Call the police? No, too much paperwork. Wake up from this bizarre dream? Preferable, but unlikely.
Summ-nP
The day started like any other. The alarm buzzed obnoxiously at 6:30 AM, shaking me from a dreamless sleep. I fumbled to silence it, already dreading the dreary commute to the office… The broadcast droned on about the usual—traffic jams, a potential storm, and the mayor’s latest scandal… until a peculiar sound interrupted my routine: a soft hum, followed by a voice behind me. "Good morning, citizen of the past!"
Table 11: Example showing the role of persona descriptions in enhancing Language Use for a Reddit author for Faithfulness to Author History. The Summ method captures rhetorical engagement, and repetition more effectively than its counterpart without persona descriptions - Summ-nP.
Author Writing Summary for Language Use
Vivid and sensory language creates immersive atmospheres. Dialogue effectively reveals character emotions and tensions. The narrative voice carries a reflective and introspective tone.
Persona Description for Language Use
Your language use is vivid and sensory, painting immersive atmospheres that draw readers into your world. Through dialogue, you reveal the emotions and tensions that simmer beneath the surface, adding authenticity and depth to your characters. Your narrative voice carries a reflective and introspective tone, inviting readers to engage with the deeper themes of your stories and reflect on their own experiences.
Writing Prompt
After moving into their dream apartment, Romi witnesses a neighbor’s tragic death, but as the days pass, she begins to suspect the act was somehow meant to shatter her own carefully constructed life.
Average Author
The apartment was a sanctuary of light. Sunlight spilled through the floor-to-ceiling windows, pooling onto the wide-plank oak floors and casting soft, dappled shadows across the walls. Romi had waited years for this—an apartment with high ceilings, a balcony overlooking the park, and an airy, open layout that made her feel as though she were floating above the city.
Summ
The apartment on the top floor had a view that stretched far beyond the horizon. Every evening, when the sun dipped below the skyline, Romi would stand by the window, tracing the fading line of light with her eyes, feeling the world shrink to just her and this slice of heaven. The move had been a leap, a daring escape from the mundane into the realm of dreams where everything seemed possible.
Summ-nP
The apartment was everything Romi had ever imagined for herself… perched high above the city with its sprawling windows that captured the sunrise like a secret whispered just for her. The floors were a deep mahogany that gleamed under the soft light… The creaking of the floors became a language of its own, a series of codes and messages that she was determined to decipher.
Table 12: Example showing that persona descriptions do not always help improve Faithfulness for expert sources like New Yorker. The Summ-nP captures vivid sensory details more effectively than the Summ when compared against the Average Story.
Table 11 shows an example for a Reddit author where the Summ method performs better that its counterpart without persona descriptions (Summ-nP) in terms of Faithfulness to Author History, evaluated using claims from the Author Writing Summary for the Language Use category. Specifically, we see that the Summ method includes rhetorical questions and repetition for emphasis that the Summ-nP method misses that make the Summ method better aligned to the Author Writing Summary.
Table 12 shows an example for a New Yorker author where the Summ method performs worse than its counterpart without persona descriptions (Summ-nP) in terms of Faithfulness to Author History, evaluated using claims from the Author Writing Summary for the Language Use category. Both the Average Author and Summ-nP better capture vivid sensory details compared to Summ, which instead emphasizes reflective and introspective tone, likely because the persona description places greater focus on this aspect.
E.2 Similarity to Author Story
E.2.1 Sheet outperforms all other methods
Writing Prompt
The world’s greatest hero has turned traitor and joined the forces of evil. Now, everyone wants to know why.
Author Story
"You were the chosen one!" the paladin screamed, pointing an accusatory finger at the prior hero. "You were to save us from evil, not join them!" The prior hero paused for a moment before nodding. "They had cookies."
The paladin was dumbfounded. "You sold away everything you were, sold out the mortal race…for *confections?*"
It is not that I am fed confections; it is that I make them, with a monster who treats me as their child, and I to her, as my mother—something I’ve never had.
It is not that I am promised riches; it is that I earn them, and am encouraged to grow by a leader who *cares* for their subjects.
Average Author
The city held its breath. The world’s former protector, now veiled in darkness, stood at the gates of ruin. "Why?" the crowd asked. But no answer came—only silence, heavy with betrayal.
His eyes, once symbols of hope, now flickered with uncertainty. A fallen hero. A world left wondering.
Sheet
In the shadow of my former self, I find clarity. "Not truth, perhaps," I admit, "but a different perspective. Perhaps the true villain is the one who blindly follows orders."
Our conversation is interrupted by a comedic spectacle—a minion, oblivious to the tension, tries to carry a stack of forms taller than himself. He stumbles, papers flying everywhere, and I can’t help but laugh.
The minion returns, this time with a clipboard. "New recruit orientation," he says sheepishly. "Welcome packet and all."
I take the clipboard, glancing over the mundane details. "Orientation, really? I thought I’d at least get a dramatic initiation ceremony."
Summary
In the heart of the once-vibrant city of Elysium, now shrouded in shadows, the air buzzed with whispers of betrayal. Streets that once echoed with laughter were silent, as newspapers screamed of the hero’s fall from grace.
"Will you stand with me, Alaric? Not as an ally of light or shadow, but as a seeker of truth?" He hesitated. Yet in the end, the promise of redemption proved irresistible. "I will," he declared. "For the sake of the city. For the sake of what we once stood for."
Table 13: Example where Sheet outperforms both Average Author and Summary for Similarity to Author Story in the Creativity category. Sheet captures the Author Story’s blend of humor and imaginative reinterpretation of the prompt, while Summary follows a more conventional hero-fall narrative and lacks the light-hearted tone.
Table 13 shows an example from the Reddit source where the Sheet method outperforms both the Average Author and Summary in terms of Similarity to the Author Story for the Creativity category. The Author Story takes a humorous and unconventional approach to the prompt, blending light-hearted banter and emotional subtext. The Sheet captures this tone through witty dialogue and imaginative reinterpretation (e.g., minion mishaps and onboarding rituals), while the Summary adopts a more serious, dramatic tone typical of heroic betrayal arcs. The Average Author response remains generic and somber, lacking the distinct creative flair that characterizes the Author Story. This example illustrates the Sheet method’s advantage in preserving the author-specific stylistic choices (unconventional re-interpretation of the prompt) beyond surface-level lexical similarity.
E.2.2 Sheet still underperforms upper-bound Oracle method
Writing Prompt
You have been in a long, lonely coma. One day, Life and Death visit your consciousness, and get into a heated debate about what should be done with you. Indifferent to your fate, you like hearing them argue like an old married couple, and decide to try and set them up with each other.
Author Story
I’ve been drifting in this vast, empty void for so long… Akin to a leaf drifting away and about in a still pond, time doesn’t really exist here, and the silence is so deep that it almost feels like it has weight… I’m trapped in some sort of elevator that not only goes up and down, but also left, right, and diagonally… And hey, if it works out, beyond recovering from my coma, I might just score a very special invite to the universe’s most unlikely wedding.
Average Author
In the quiet echo of my mind, where only the soft hum of forgotten dreams lingered… It was like waking up, but not quite; it was more like tuning into a radio station I hadn’t realized existed… It was like listening to an old married couple who’d had the same argument for centuries… I imagined them, Life with her radiant glow, hands on her hips, eyes sparkling with the passion of a thousand dawns, and Death, calm and composed, a gentle shadow with a hint of a smile playing on his lips.
Sheet
Consciousness is a peculiar thing. It floats somewhere between the tangible and the illusory, like a whisper that hangs in the air, just out of reach. In the coma world, where I have resided for what feels like an eternity, time has no meaning… "Wouldn’t it be easier if you both just worked together?" I muse, my thoughts a ripple in the ether.
Life laughs, a sound like tinkling bells. "Oh, wouldn’t that be something? But we are not meant to intertwine. We serve different purposes."… And in this place of dreams and whispers, I choose to live, fully and freely, in whatever form that may take.
Oracle
Floating in the vast, ethereal void of my own consciousness was akin to being a leaf drifting in a still pond—untouched by time, unanchored by space, and entirely alone. "Have you ever considered that maybe all this arguing is just… romantic tension?"
Life blinked, momentarily at a loss for words. Death’s lips curled into a faint smile, his eyes twinkling with amusement. "Romantic tension?"
Table 14: Example for a Reddit author where Sheet underperforms Oracle in capturing the Author Story’s lively and light-hearted tone and dialogue and instead leans toward introspection and philosophical reflection.
Table 14 shows an example for a Reddit author where the Sheet method loses to the Average Author, while the Oracle method achieves the best performance for the Language Use category. The Sheet’s primary limitation is its lack of lively, humorous dialogue exchanges present in both the Author Story and the Oracle. Instead, it leans toward introspection and philosophical reflection. In contrast, the Average Author includes light-hearted descriptions that bring it closer to the Author Story, though it still falls short of the Oracle, which captures the tone and dialogue more effectively.
E.2.3 Effect of Persona on Similarity to Author Story